OpenStudy (anonymous):

What is the exact value of sin(2 theta) if cos theta =3/5 and is in quadrant IV?

OpenStudy (anonymous):

quadrant IV is quadrant 4 which is the one in the bottom right hand corner

OpenStudy (michele_laino):

we have: \[\Large \sin \theta = - \sqrt {1 - {{\left( {\cos \theta } \right)}^2}} = - \sqrt {1 - \frac{9}{{25}}} = ..?\]

OpenStudy (michele_laino):

the minus sign takes account that we are in the IV quadrant, where sin x is negative. So we have to apply this formula: \[\Large \sin \left( {2\theta } \right) = 2\sin \theta \times \cos \theta = ...?\]